Generally speaking, would you say that most people can be trusted? A random sample of n1 = 255 people in Chicago ages 18-25 showed that r1 = 43 said yes. Another random sample of n2 = 270 people in Chicago ages 35-45 showed that r2 = 65 said yes. Does this indicate that the population proportion of trusting people in Chicago is higher for the older group? Use α = 0.05.

a)What is the value of the sample test statistic? (Test the difference p1p2. Do not use rounded values. Round your final answer to two decimal places.)

b) Find (or estimate) the P-value. (Round your answer to four decimal places.
Answer #1

P1 :- Proportion of people in Chicago ages 18-25

P2 :- Proportion of people in Chicago ages 35-45

To Test :-

H0 :- P1 = p2

H1 :- P1 < P2



Test Statistic :-

is the pooled estimate of the proportion P
= ( x1 + x2) / ( n1 + n2)
= ( 43 + 65 ) / ( 255 + 270 )
= 0.2057

Z = -2.04


Test Criteria :-
Reject null hypothesis if

= -2.04 < -1.64, hence we reject the null hypothesis
Conclusion :- We Reject H0


Decision based on P value
P value = P ( Z < -2.04 )
P value = 0.0207
Reject null hypothesis if P value <
Since P value = 0.0207 < 0.05, hence we reject the null hypothesis
Conclusion :- We Reject H0

There is sufficient evidence to support the claim that the population proportion of trusting people in Chicago is higher for the older group at 5% level of significance.